Time and work question

A alone can do a piece of work in 6 days and B alone in 8 days. A and B undertook to do it for Rs. 3200. With the help of C, they completed the work in 3 days. How much is to be paid to C?

A.    Rs. 375    B.    Rs. 400
C.    Rs. 600    D.    Rs. 800

Grade:upto college level

1 Answers

Amit Saxena
10 years ago

Answer: Option B

Explanation:

C''s 1 day''s work = 1 - 1 + 1 = 1 - 7 = 1 .
3 6 8 3 24 24
A''s wages : B''s wages : C''s wages = 1 : 1 : 1 = 4 : 3 : 1.
6 8 24
C''s share (for 3 days) = Rs. 3 x 1 x 3200 = Rs. 400.
24